# CNRS &\& Sorbonne Université, 4 Place Jussieu, Paris, France

### v2

@@ 2.1 Vector portal: dark photon or kinetic-mixing model
As discussed above, we obtain the DM-induced annihilation spectra of $e^{\pm}$ and photons using HAZMA. A few examples of the $e^{\pm}$ spectrum $\left.\frac{dN_{e}}{dE_{e}}\right|_{\rm tot}$ for different DM masses ($m_{\rm DM}=30,300$ and 700 MeV) are shown in the left panel of fig. 1. Here $\left.\frac{dN_{e}}{dE_{e}}\right|_{\rm tot}=\sum_{F}B_{F}\,\frac{dN^{F}_{e}}{dE_{e}}$ is the total spectrum as a sum of the contribution from each (kinematically possible) annihilation channel $F$ with the corresponding branching fraction $B_{F}$. The spectra consist of two main components. A sharp peak component, coming from the mediator ($V^{*}$) decaying into $e^{+}\,e^{-}$, dominates for lower $m_{\rm DM}$’s and at higher energies for larger $m_{\rm DM}$’s. There is an additional smooth component (relevant for comparatively larger $m_{\rm DM}$’s), coming from mediator decays into muons and combinations of light mesons, which dominates the low energy part of the spectrum. See ref. Coogan:2022cdd for details. Note that in our parameter setup ($m_{V}=3\,m_{\rm DM}$) the spectra $\left.\frac{dN}{dE}\right|_{\rm tot}$ depend only on $m_{\rm DM}$ (or $m_{V}$), but not on the choice of $g_{V{\rm DM}}$ or $\epsilon$, since the latter parameters enter only in determining the strength of $V^{*}$ production and decay. In the opposite regime ($m_{V}<m_{\rm DM}$), the competing annihilation channel ${\rm DM}\,{\rm DM}\to VV\to f\bar{f}f^{\prime}\bar{f}^{\prime}$ (‘Secluded Annihilation’) opens up, producing different spectra, and the interplay among the two processes (hence the total spectra) depends on all parameters.
@@ 2.2 Scalar portal: higgs-portal model
A few examples of the total $e^{\pm}$ spectrum $\left(\left.\frac{dN_{e}}{dE_{e}}\right|_{\rm tot}\right)$ in this model are shown in the right panel of fig. 1 for different DM masses ($m_{\rm DM}=30,300$ and 700 MeV). The spectra consist of a box component coming from the on-shell mediator $S$ decaying into $e^{+}\,e^{-}$ (which dominates at lower DM masses for $m_{\rm DM}$ above the electron threshold, 2 MeV), and a smooth component coming from $S$ decaying into muons and pions (which is dominating at higher DM masses). See Coogan:2019qpu ; Coogan:2021rez ; Coogan:2021sjs for details. Note that in our parameter regime ($m_{S}=m_{\rm DM}/2$) the spectra $\left.\frac{dN}{dE}\right|_{\rm tot}$ for $e^{\pm}$ or photon depends only on $m_{\rm DM}$ (or $m_{S}$), but not on the choice of $g_{S{\rm DM}}$ or $\sin\vartheta$ (at least for not-too-small values of these parameters, otherwise the competing annihilation channel DM DM $\to S\to f\bar{f}$ becomes relevant and, since it depends $g_{S{\rm DM}}$ and $\sin\vartheta$, the total spectra are modified).
@@ 3 Photon signals from sub-GeV DM annihilation in the inner Galaxy
with $r$ as the radial distance from the galactic center (GC). For the parameters $\rho_{0}$ and $r_{s}$ we consider the values corresponding to the fitted NFW profile parameters (the central values) obtained in 2019JCAP…10..037D for the baryonic model B2. However, for comparison we will also consider other choices for the DM profile, see App. B.
@@ 6.4.2 CMB constraints
In each $\langle\sigma v\rangle-m_{\rm DM}$ plane we show the constraint from Planck CMB observation Planck:2018vyg with a gray dash-dotted line. This is estimated for the two considered sub-GeV DM models (where DM is a Dirac fermion with particle-antiparticle symmetry) following Coogan:2019qpu with the efficiency factors for $e^{\pm}$ and $\gamma$ from Slatyer:2015jla . The CMB constraint for the kinetic-mixing model is comparatively much stronger (compared to the thermal line) since here the DM annihilation is $s$-wave. On the other hand, this constraint in the case of the higgs-portal model is much weaker since here the annihilation is $p$-wave. In this case, the DM velocity at recombination which is required to translate the CMB constraint on $\langle\sigma v\rangle$ to its present-day value is estimated considering an optimistic choice for the kinetic decoupling temperature of $x_{kd}=10^{-6}$ (increasing which should relax the CMB constraint) Essig:2013goa .
@@ 6.4.3 Terrestrial constraints
For the higgs-portal model with $m_{S}<m_{\rm DM}$, the DM pair-annihilation cross-section depends only on the coupling $g_{S{\rm DM}}$, but not on the the mixing parameter $\sin\vartheta$ between the scalar mediator and the SM higgs. The only requirement is that there is some value of $\sin\vartheta$ consistent with the terrestrial observations, and that it is large enough that the decay length of the mediator is below the parsec scale Coogan:2021sjs . Hence, the complementarity between the indirect detection constraints and those from the terrestrial searches Krnjaic:2015mbs can not be shown in the $\langle\sigma v\rangle-m_{\rm DM}$ plane. For this model, we compare our results with the cosmological constraints (CMB and relic abundance). In addition, like in Coogan:2021sjs , we show (in fig. 5) some contours of constant $g_{S{\rm DM}}$ to get an idea of reasonable values of $\langle\sigma v\rangle$. For example, the existing photon and cosmic-ray observations together are able to constrain the values of $g_{S{\rm DM}}$ between $\sim 0.1$ – a few $\times 10^{-3}$ for a DM mass up to a few hundreds of MeV. The near-future Cosi will be able to improve this situation further.
@@ Acknowledgments
We thank Peter Reimitz for useful discussions and collaboration in the initial stages of this work. We also thank Pierre Salati for providing us with the Voyager-1 data and Lukas Simon for a useful discussion. M.C and A.K. acknowledge the hospitality of the Institut d’Astrophysique de Paris (Iap) where part of this work was done. M.C. also acknowledges the hospitality of the Flatiron Institute (Simons Foundation) and of the Physics Department of New York University.
@@ Appendix B Dependence on the DM profile
In fig. 8 we illustrate the effect of considering different DM profiles on the projected sensitivities of Cosi, for the two considered DM models. In each case, the brown solid line corresponds to the projection obtained with the NFW profile of eq. 3 and is same as the one shown in fig. 6. The associated brown band indicates the variation due to different choices for the DM profile. The upper edge of the band corresponds to the truncated Isothermal (cored) profile with the corresponding parameters tabulated in Cirelli:2024ssz (also used in Cirelli:2025qxx ). On the other hand, the lower edge of the band corresponds to the Einasto profile from 2019JCAP…10..037D (for the baryonic model B2) with the profile parameters adjusted within their $1\sigma$ uncertainty to maximize the $J$-factor (like in Coogan:2021sjs ). This figure shows that, depending on the choice of DM profile, the indirect detection constraints on the DM models presented in fig. 6 can vary by a factor of a few in both directions.

### v3

@@ 2.1 Vector portal: dark photon or kinetic-mixing model
As discussed above, we obtain the DM-induced annihilation spectra of $e^{\pm}$ and photons using HAZMA. A few examples of the $e^{\pm}$ spectrum $\left.\frac{dN_{e}}{dE_{e}}\right|_{\rm tot}$ for different DM masses ($m_{\rm DM}=30,300$ and 700 MeV) are shown in the left panel of fig. 1. Here $\left.\frac{dN_{e}}{dE_{e}}\right|_{\rm tot}=\sum_{F}B_{F}\,\frac{dN^{F}_{e}}{dE_{e}}$ is the total spectrum as a sum of the contribution from each (kinematically possible) annihilation channel $F$ with the corresponding branching fraction $B_{F}$. The spectra consist of two main components. A sharp peak component, coming from the mediator ($V^{*}$) decaying into $e^{+}\,e^{-}$, dominates for lower $m_{\rm DM}$’s and at higher energies for larger $m_{\rm DM}$’s. There is an additional smooth component (relevant for comparatively larger $m_{\rm DM}$’s), coming from mediator decays into muons and combinations of light mesons, which dominates the low energy part of the spectrum. See ref. Coogan:2022cdd for details. Note that in our parameter setup ($m_{V}=3\,m_{\rm DM}$) the spectra $\left.\frac{dN}{dE}\right|_{\rm tot}$ depend only on $m_{\rm DM}$ (or $m_{V}$), but not on the choice of $g_{V{\rm DM}}$ or $\epsilon$  [^3] In general, $\left.\frac{dN}{dE}\right|_{\rm tot}=\sum_{F}B_{F}\frac{dN^{F}}{dE}(m_{\rm DM})=\sum_{F}\frac{\langle\sigma v\rangle_{F}}{\langle\sigma v\rangle_{\rm tot}}\frac{dN^{F}}{dE}(m_{\rm DM})$. Here $\frac{dN^{F}}{dE}$ is the $\gamma$ or $e^{+}$ spectrum produced in the SM channel $F$ and depends only on the outgoing energy in the channel $F$ (i.e., on the DM mass) but not on the DM couplings. In the case with $m_{V}=3\,m_{\rm DM}$, the only kinematically allowed annihilation channels are DM DM $\to V^{*}\to f\bar{f}$, for which $\langle\sigma v\rangle_{F}\propto(\epsilon\,g_{V{\rm DM}})^{2}$ (see Eq. 1), and hence $\langle\sigma v\rangle_{\rm tot}=\sum_{F}\langle\sigma v\rangle_{F}\propto(\epsilon\,g_{V{\rm DM}})^{2}$. As a result, the spectra $\left.\frac{dN}{dE}\right|_{\rm tot}$ are independent on $g_{V{\rm DM}}$ or $\epsilon$., since the latter parameters enter only in determining the strength of $V^{*}$ production and decay. In the opposite regime ($m_{V}<m_{\rm DM}$), the competing annihilation channel ${\rm DM}\,{\rm DM}\to VV\to f\bar{f}f^{\prime}\bar{f}^{\prime}$ (‘Secluded Annihilation’) opens up, producing different spectra, and the interplay among the two processes (hence the total spectra) depends on all parameters.
@@ 2.2 Scalar portal: higgs-portal model
A few examples of the total $e^{\pm}$ spectrum $\left(\left.\frac{dN_{e}}{dE_{e}}\right|_{\rm tot}\right)$ in this model are shown in the right panel of fig. 1 for different DM masses ($m_{\rm DM}=30,300$ and 700 MeV). The spectra consist of a box component coming from the on-shell mediator $S$ decaying into $e^{+}\,e^{-}$ (which dominates at lower DM masses for $m_{\rm DM}$ above the electron threshold, 2 MeV), and a smooth component coming from $S$ decaying into muons and pions (which is dominating at higher DM masses). See Coogan:2019qpu ; Coogan:2021rez ; Coogan:2021sjs for details. Note that in our parameter regime ($m_{S}=m_{\rm DM}/2$) the spectra $\left.\frac{dN}{dE}\right|_{\rm tot}$ for $e^{\pm}$ or photon depends only on $m_{\rm DM}$ (or $m_{S}$), but not on the choice of $g_{S{\rm DM}}$ or $\sin\vartheta$ (at least for not-too-extreme values of these parameters, otherwise the competing annihilation channel DM DM $\to S\to f\bar{f}$ becomes relevant and, since it depends $g_{S{\rm DM}}$ and $\sin\vartheta$, the total spectra are modified)  [^4] As mentioned in sec. 2.1, $\left.\frac{dN}{dE}\right|_{\rm tot}=\sum_{F}\frac{\langle\sigma v\rangle_{F}}{\langle\sigma v\rangle_{\rm tot}}\frac{dN^{F}}{dE}(m_{\rm DM})$. As long as only ${\rm DM}\,{\rm DM}\to VV\to f\bar{f}f^{\prime}\bar{f}^{\prime}$ annihilation dominates, $\langle\sigma v\rangle_{F}$ and thus $\langle\sigma v\rangle_{\rm tot}$ $(=\sum_{F}\langle\sigma v\rangle_{F})$ are proportional to $g^{4}_{S{\rm DM}}$ (Eq. 2). Hence, $\left.\frac{dN}{dE}\right|_{\rm tot}$ is independent on $g_{S{\rm DM}}$ or $\sin\vartheta$..
@@ 3 Photon signals from sub-GeV DM annihilation in the inner Galaxy
with $r$ as the radial distance from the galactic center (GC). For the parameters $\rho_{0}$ and $r_{s}$ we consider the values $\rho_{0}=0.839\,{\rm GeV\,cm^{-3}}$ and $r_{s}=11$ kpc, which correspond to the fitted NFW profile parameters (the central values) obtained in 2019JCAP…10..037D for the baryonic model B2. However, for comparison we will also consider other choices for the DM profile, see App. B.
@@ 6.4.2 CMB constraints
In each $\langle\sigma v\rangle-m_{\rm DM}$ plane we show the constraint from Planck CMB observation Planck:2018vyg with a gray dash-dotted line. This is estimated for the two considered sub-GeV DM models (where DM is a Dirac fermion with particle-antiparticle symmetry) following Coogan:2019qpu with the efficiency factors for $e^{\pm}$ and $\gamma$ from Slatyer:2015jla . The CMB constraint for the kinetic-mixing model is comparatively much stronger (compared to the thermal line) since here the DM annihilation is $s$-wave. On the other hand, this constraint in the case of the higgs-portal model is much weaker since here the annihilation is $p$-wave. In this case, the DM velocity at recombination which is required to translate the CMB constraint on $\langle\sigma v\rangle$ to its present-day value is estimated considering an optimistic choice for the kinetic decoupling temperature of $x_{kd}=10^{-6}$ (increasing which should relax the CMB constraint) Essig:2013goa . Note here that, in the cases in which the vector portal model also features $p$-wave annihilations (e.g. if the DM is a complex scalar or a Majorana fermion), the CMB constraints are equally relaxed and the corresponding thermal relic lines go down. Instead the indirect detection limits obtained in our analysis are sensitive only to DM annihilation at the present time and therefore are quite independent of the $s$-wave/$p$-wave assumption if we consider, as usually done, a constant DM speed in the galactic halo.
@@ 6.4.3 Terrestrial constraints
For the higgs-portal model with $m_{S}<m_{\rm DM}$, the DM pair-annihilation cross-section depends only on the coupling $g_{S{\rm DM}}$, but not on the mixing parameter $\sin\vartheta$ between the scalar mediator and the SM higgs. The only requirement is that there is some value of $\sin\vartheta$ consistent with the terrestrial observations, and that it is large enough that the decay length of the mediator is below the parsec scale Coogan:2021sjs . Hence, the complementarity between the indirect detection constraints and those from the terrestrial searches Krnjaic:2015mbs can not be shown in the $\langle\sigma v\rangle-m_{\rm DM}$ plane. For this model, we compare our results with the cosmological constraints (CMB and relic abundance). In addition, like in Coogan:2021sjs , we show (in fig. 5) some contours of constant $g_{S{\rm DM}}$ to get an idea of reasonable values of $\langle\sigma v\rangle$. For example, the existing photon and cosmic-ray observations together are able to constrain the values of $g_{S{\rm DM}}$ between $\sim 0.1$ – a few $\times 10^{-3}$ for a DM mass up to a few hundreds of MeV. The near-future Cosi will be able to improve this situation further.
6.4.4 SN1987A constraints
For the kinetic-mixing model, we also compare our DM indirect detection results with the Supernova 1987A (SN1987A) constraints on the same model from Chang:2018rso . Considering the fiducial SN model and a choice for the coupling parameter $\alpha_{D}=0.5$ (like in the case of the terrestrial constraints discussed above), we map the region of the parameter space excluded by SN1987A from the $y-m_{\rm DM}$ plane (presented in the right panel of figure 6 of Chang:2018rso ) into our $\langle\sigma v\rangle-m_{\rm DM}$ plane (following the same prescription discussed above in sec. 6.4.3). We find that this excludes a region $10^{-37}\,{\rm cm^{3}\,s^{-1}}\lesssim\langle\sigma v\rangle\lesssim 10^{-31}\,{\rm cm^{3}\,s^{-1}}$ (which falls below the level shown in fig. 5 or 6) for $m_{\rm DM}$ in the range 1 – $\sim 100$ MeV, leaving the other parts of the parameter space, in particular those to be probed by upcoming DM indirect detection searches, allowed.
@@ Acknowledgments
We thank Peter Reimitz for useful discussions and collaboration in the initial stages of this work. We thank Pierre Salati for providing us with the Voyager-1 data. We also thank Lukas Simon and Andrea Caputo for useful discussions. M.C and A.K. acknowledge the hospitality of the Institut d’Astrophysique de Paris (Iap) where part of this work was done. M.C. also acknowledges the hospitality of the Flatiron Institute (Simons Foundation) and of the Physics Department of New York University.
@@ Appendix B Dependence on the DM profile
In fig. 8 we illustrate the effect of considering different DM profiles on the projected sensitivities of Cosi, for the two considered DM models. In each case, the brown solid line corresponds to the projection obtained with the NFW profile of eq. 3 and is same as the one shown in fig. 6. The associated brown band indicates the variation due to different choices for the DM profile. The upper edge of the band corresponds to the truncated Isothermal (cored) profile $\rho^{\rm Iso}_{\rm DM}(r)={\rho_{0}}/\left(1+\left(\frac{r}{r_{s}}\right)^{2}\right)$, with the corresponding parameters $\rho_{0}=2.112\,{\rm GeV\,cm^{-3}}$ and $r_{s}=4$ kpc tabulated in Cirelli:2024ssz (also used in Cirelli:2025qxx ). On the other hand, the lower edge of the band corresponds to the Einasto profile $\rho^{\rm Ein}_{\rm DM}(r)=\rho_{0}\,{\rm exp}\left\{-\frac{2}{\alpha}\left(\left(\frac{r}{r_{s}}\right)^{\alpha}-1\right)\right\}$ from 2019JCAP…10..037D (corresponding the baryonic model B2) with the profile parameters adjusted within their $1\sigma$ uncertainty to maximize the $J$-factor (like in Coogan:2021sjs ). This corresponds to $\alpha\simeq 0.1$, $r_{s}\simeq 6.5$ kpc and $\rho_{\rm DM}(r_{\odot})\simeq 0.4\,{\rm GeV\,cm^{-3}}$. Fig. 8 shows that, depending on the choice of DM profile, the indirect detection constraints on the DM models presented in fig. 6 can vary by a factor of a few in both directions.